1. * + 4x + 8x = e, x(0) = x(0) = 0.

Use the Laplace transform to solve initial value problems
Answer #1

(L st 1 At Stl O *+421 +82 = ēt, 460) = x0) = 0 Take Laplace Transform on both sides Hat]+446_!] +84(] =l[et] sixecs) –S260)5+1 6+45+8) I 5(541) + 35-315 St4578 Y(5) = .S+3 5+4548 = (59% 5(sti) - ( 3+2+1 5+2+1 ($+2²+4 XLS) - T 5(SHU) $ ( &#2374 + LL